Aswhi7424:

Suppose the sample space for a continuous random variable is to 80. the area under the density curve for the variable from 0 to 5 is 0.0625what is the area under the density curve from 5 to 80?

badgirlhope:

What do you think it is?

Aswhi7424:

I have no clue how to even set it up.

supie:

You have to solve \[p(x>5)=1-p(x<5)=1-0.0625\] right? Also, choices would help.